# Evaluating Agents using Social Choice Theory

### v2

@@ 3.1.1 Probabilistic Voting: Maximal Lotteries
1 Let $W\leftarrow()$ be an empty sequence for round $t\in\{0,1,2,\cdots,m-1\}$ do
@@
4      $x_{t}\leftarrow\textsc{MaximalLotteries}(A_{t})$ $w_{t}\leftarrow\{a\in A_{t}:x_{t}(a)>0\}$ Append $(x_{t},w_{t})$ to the end of $W$ $A_{t+1}=A_{t}-w_{t}$ Remove the winners $w_{t}$ from all the votes.
return $W$
@@
We propose an iterative version of the maximal lotteries depicted in Algorithm 1, which repeatedly removes winners and solves the margin sub-games among the remaining agents. Iterative ML (IML) retains the strong properties of ML (Brandl and Brandt, 2020). Additionally, as shown in our evaluation, it can identify game-theoretic cycles among subsets of agents. Details about this iterative approach can be found in the appendix.
@@ 4.1 General Reinforcement Learning Agents
Observation 4: Nash averaging requires score normalization. Recall that Nash averaging (Balduzzi et al., 2018) finds an equilibrium in the Agent-versus-Task matrix induced by (Hessel et al., 2017, Table 5). If we were to run Nash averaging on raw scores, the task player would place full weight on one task (skiing) in its equilibrium because the rewards are large and negative ($<-10000$) and outside the bounds of the other games, across all agents. To avoid this pathology, our Nash Averaging implementation
first normalizes scores into the range of $[0,1]$ as in previous work (Balduzzi et al., 2018). In contrast, VasE compares agents’ raw scores only within the same task, without requiring normalization between different tasks.
[…]
@@ 4.2 Ranking Large Language Models as Agents
[…]
@@ C.2.9 Maximal Lotteries
1 Let $W\leftarrow()$ be an empty sequence for round $t\in\{0,1,2,\cdots,m-1\}$ do
@@
4      $x_{t}\leftarrow\textsc{MaximalLotteries}(A_{t})$ $w_{t}\leftarrow\{a\in A_{t}:x_{t}(a)>0\}$ Append $(x_{t},w_{t})$ to the end of $W$ $A_{t+1}=A_{t}-w_{t}$ Remove the winners $w_{t}$ from all the votes.
return $W$
@@
A reasonable question to ask is if this iterative version of maximal lotteries preserves the properties of the original voting scheme. The answer is yes. In particular, by leveraging agenda consistency (Definition 5) or equivalently Sen’s $\alpha$-contraction and $\gamma$-expansion properties (Sen, 1971, 1977), one can “safely” remove alternatives that are not in the support of the lottery returned by the maximal lottery as they have no effect on the outcome. Issues, like the lack of independence of irrelevant alternatives that drive many impossibility results when working with deterministic voting rules, no longer apply to maximal lotteries (Brandl and Brandt, 2020).
@@ 13 RWKV-4-Raven-14B 499
[…]

### v3

@@ 3.1.1 Probabilistic Voting: Maximal Lotteries
1 Let $W\leftarrow()$ be an empty sequence
2 for round $t\in\{0,1,2,\cdots,m-1\}$ do
@@ 1 Let $W\leftarrow()$ be an empty sequence
5
6      $x_{t}\leftarrow\textsc{MaximalLotteries}(A_{t})$
7       $w_{t}\leftarrow\{a\in A_{t}:x_{t}(a)>0\}$
8       Append $(x_{t},w_{t})$ to the end of $W$
9       $A_{t+1}=A_{t}-w_{t}$
10       Remove the winners $w_{t}$ from all the votes.
11
12return $W$
@@ 9       $A_{t+1}=A_{t}-w_{t}$
We propose Iterative Maximal Lotteries (IML), depicted in Algorithm 1, which repeatedly removes winners and solves the margin sub-games among the remaining agents. As shown in our evaluation, IML can identify game-theoretic cycles among subsets of agents. Details about this iterative approach can be found in the appendix.
@@ 4.1 General Reinforcement Learning Agents
Observation 4: Nash averaging requires score normalization. Recall that Nash averaging (Balduzzi et al., 2018) finds an equilibrium in the Agent-versus-Task matrix induced by (Hessel et al., 2017, Table 5). If we were to run Nash averaging on raw scores, the task player would place full weight on one task (skiing) in its equilibrium because the rewards are large and negative ($<-10000$) and outside the bounds of the other games, across all agents. To avoid this pathology, our Nash Averaging implementation first normalizes scores into the range of $[0,1]$ as in previous work (Balduzzi et al., 2018). In contrast, VasE compares agents’ raw scores only within the same task, without requiring normalization between different tasks.
Observation 5: Nash averaging has adversarial task selection in ALE. With normalization, Nash averaging finds a four-way tie for first place {a3c, dist-dqn, rainbow, duel}, with a task distribution of (assault: 0.36,boxing: 0.29,breakout: 0.26,venture: $(\texttt{assault}\text{:}\leavevmode\nobreak\ 0.36,\leavevmode\nobreak\ \allowbreak\texttt{boxing}\text{:}\leavevmode\nobreak\ 0.29,\leavevmode\nobreak\ \allowbreak\texttt{breakout}\text{:}\leavevmode\nobreak\ 0.26,\leavevmode\nobreak\ \allowbreak\texttt{venture}\text{:}\leavevmode\nobreak\ 0.09)$. The consequence of the approach is clear: first, Nash equilibrium guarantees that the top agents in the support have the same value ($0.3891$ in this case) making them indistinguishable under this metric. Second, due to adversarial task selection, only four (7.4% of) tasks are used to rank agents.
@@ 4.2 Ranking Large Language Models as Agents
Observation 8: Nash averaging suffers from adversarial task selection in AgentBench. We run Nash averaging on AgentBench in a similar way as on ALE, by normalizing the scores into $[0,1]$. As in ALE, Nash averaging finds a low-support equilibrium, with a task distribution of (HouseHolding: 0.135,WebShopping: $(\texttt{HouseHolding}\text{:}\leavevmode\nobreak\ 0.135,\allowbreak\texttt{WebShopping}\text{:}\leavevmode\nobreak\ 0.865)$. Again, a large portion ($\frac{6}{8}=75\%$) of the tasks are not used in ranking the agents. The task player places most weight on the only task that gpt-4 does not place 1st. Under this distribution, there is a two-way tie for top rank between gpt-4 and gpt-3.5-turbo with a Nash average value of 0.889.
@@ 5 Conclusion and Future Work
5.1 Acknowledgments
We would like to thank Manfred Diaz for continued discussions on evaluation methodologies and Felix Brandt for his feedback on earlier versions of this paper.
@@ C.2.9 Maximal Lotteries
1 Let $W\leftarrow()$ be an empty sequence
2 for round $t\in\{0,1,2,\cdots,m-1\}$ do
@@ 1 Let $W\leftarrow()$ be an empty sequence
5
6      $x_{t}\leftarrow\textsc{MaximalLotteries}(A_{t})$
7       $w_{t}\leftarrow\{a\in A_{t}:x_{t}(a)>0\}$
8       Append $(x_{t},w_{t})$ to the end of $W$
9       $A_{t+1}=A_{t}-w_{t}$
10       Remove the winners $w_{t}$ from all the votes.
11
12return $W$
@@ 13 RWKV-4-Raven-14B 499
In the template shown in Figure 15, $<$example full texti$>$ and $<$example summary i$>$ are manually created full text descriptions of a destination and a manually created summary text for that destination. The prompts to be evaluated are injected into the above template in the $<$specific prompt$>$ location. In our evaluation we consider the following prompts (clearly many other prompts can be evaluated in a similar way).